Emily is 26 years older than her daughter Maria In 10 years emily will be three times as old as maria how old is each at present?
Answer by Maths68(1474)

Let
Present age of Emily = e
Present age of Maria = m
Given
Emily is 26 years older than her daughter Maria means if add 26 to the present age of Maria both Emily and Maria will be of equal age
e=m+26.............(1)
In 10 years means we have to add 10 in the present age of Maria and Emily
e+10=m+10
then
Emily will be three times as old as maria
e+10=3(m+10)
e+10=3m+30
e=3m+30-10
e=3m+20...............(2)
Put the value of e from (1) to (2)
m+26=3m+20
m=3m+20-26
m-3m=-6
-2m=-6
Divide above equation by -2
-2m/-2=-6/-2
m=3
Put the value of m in (1)
e=m+26.............(1)
e=3+26
e=29
Present age of Emily = e = 29 years old
Present age of Maria = m = 3 years old